naphthalene, a white solid used to make mothballs,has a vapour pressure of 0.10 mm hg at 27degree celcius. hence,Kp and Kc for the equilibrium are : C 10H8 (s) eqiulibrium with C10H8(g).

Grade:Upto college level

3 Answers

Vikash Kumar
10 years ago
Kp = Kc (RT) ?n Where is number of moles of gaseous products – number of moles of gaseous reactants in the balanced chemical reaction.
so here there is only one gas molecule in product so Kp = .1 mm = 1.316 * 10-4 atm. (since 760mm=1 atm)
 ?n=1
so  Kc = Kp /RT = (1.316 * 10-4 )/(.0821 * 300) = 5.34 * 10-6 
Thanks & Regards
Vikash Kumar, 
askIITians Faculty
B.Tech, IIT Delhi
hemanth
10 years ago

p = Kc (RT) ?n Where is number of moles of gaseous products – number of moles of gaseous reactants in the balanced chemical reaction.
so here there is only one gas molecule in product so Kp = .1 mm = 1.316 * 10-4 atm. (since 760mm=1 atm)
 ?n=1
so  Kc = Kp /RT = (1.316 * 10-4 )/(.0821 * 300) = 5.34 * 10-6 
Thanks & Regards
Vikash Kumar, 
askIITians Faculty
B.Tech, IIT Delhi
Raunaq Mehta
7 years ago
p = Kc (RT) ?n Where is number of moles of gaseous products – number of moles of gaseous reactants in the balanced chemical reaction.so here there is only one gas molecule in product so Kp = .1 mm = 1.316 * 10-4 atm. (since 760mm=1 atm) ?n=1so Kc = Kp /RT = (1.316 * 10-4 )/(.0821 * 300) = 5.34 * 10-6